# The direct linearization scheme with the Lamé function: The KP equation and reductions

### v2

@@ 1 Introduction
A quite recent progress of the DL approach is the establishment of its elliptic scheme involving infinite matrix structure and ‘discrete’ Lamé-type PWFs [31]. The Lamé function is a solution of the Lamé equation (known as the Schrödinger equation with an elliptic potential) $y^{\prime\prime}+(A+B\wp(x))y=0$ when taking $A=-\wp(k)$ and $B=-2$. The solution reads $y=\frac{\sigma(x+k)}{\sigma(x)\sigma(k)}e^{-\zeta(k)x}$, where $\sigma,\zeta$ (and Weierstrass-$\wp$) are the Weierstrass functions. For the KdV equation, there are exact solutions involving the Lamé function, which emerged in [18] and [39] 1970s. In 2010, Nijhoff and Atkinson [25] employed a discrete analogue of the Lamé function and constructed solutions for the Adler-Bobenko-Suris [1] (except Q4 equation) lattice equations. The obtained solutions are presented in terms of the elliptic Cauchy matrix and called ‘elliptic solitons’ to distinguish from the algebro-geometric (finite-gap) solutions (cf.[5]). Later, their treatment was extended to the lattice KP equations [42]. The elliptic DL scheme established in [31] is based on the infinite matrix structures. The concept of ‘elliptic $N$th root of unity’ was introduced to define discrete dispersion relations as well as make reductions from the lattice KP system to the lower dimensional equations, namely, the discrete KdV and Boussinesq equations. As for the elliptic solitons for the continuous equations, a recent progress is due to [20], which established a framework of bilinear approach that involves the Lamé function, including formulae of $\tau$ functions, vertex operators and bilinear identities for the KP hierarchy and their reductions. Periodic degenerations (i.e. trigonometric/hyperbolic and rational forms) of the elliptic matter were also discussed in [20]. Later, the study of [20] was extended to discrete case in [21]. More recently, the consideration of such type of elliptic soliton solutions from Sato’s approach was developed by Kakei [17], where the Darboux transformation based on pseudo-differential operators and the Lamé function was established getting elliptic $N$-soliton solutions for the whole KP hierarchy and the KdV hierarchy. The elliptic solitons are also related to a recent work by Nakayashiki [23].
@@
In this paper, we will extend the Fokas-Ablowitz DL scheme to the elliptic soliton case and formulate a mechanism to obtain non-singular real-valued elliptic soliton solutions. First, for the KP equation, we will present an integral equation involving elliptic PWFs and Lamé function, using which we can define a wave function and solution for the KP equation. This DL scheme will be confirmed by checking the wave function and solution satisfy the KP Lax pair. Real-valuedness of the Weierstrass functions will be discussed (see Appendix A). In light of this, we are able to construct a Marchenko equation for elliptic soliton solutions of the KP equation. By taking special measure in the integral equation, we may obtain explicit expressions for the elliptic solitons (as well as $\tau$ functions) of the KP equation. Some solutions will be illustrated. In principle, employing the notion of elliptic $N$th root of unity introduced in [31], we can reduce the elliptic DL scheme, Marchenko equation and solutions of the KP equation to those of the KdV equation and the Boussinesq equation. However, in practice, there are many details to be fixed, which will be elaborated in the paper.
@@
We organize the paper as follows. In section 2, we set up the elliptic DL scheme for the KP equation, construct its Marchenko equation and derive formulae of the elliptic $N$-soliton solution and $\tau$-function. Then, in section 3, we elaborate reduction details so that the elliptic DL schemes, Marchenko equations and elliptic solitons can be constructed for the KdV equation and Boussinesq equation. Conclusions are given in section 4. There are two Appendix sections. One is devoted to the real-valuedness of the Weierstrass functions. The other consists of the analysis about elliptic cube roots.
@@ 2.1 Elliptic DL scheme
and the KP equation arises from the compatibility $[P,M]=PM-MP=0$. When $u(x,y,t)=u_{0}(x)$, the above Lax pair admits a special solution
@@
$\wp(z+2w_{i})=\wp(z),~{}~{}i=1,2,$
where $w_{1}$ and $w_{2}$ are two half periods. $\zeta(z)$ and $\sigma(z)$ are odd and quasi-periodic with respect to $w_{i}$, in the sense that
$\zeta(z+2w_{i})=\zeta(z)+2\zeta(w_{i}),$ (2.7a)
$\sigma(z+2w_{i})=-\sigma(z)e^{2\zeta(w_{i})(z+w_{i})},\quad i=1,2.$ (2.7b)
When $k\neq w_{i}$, $\Psi_{x}(k)$ and $\Psi_{x}(-k)$ are linearly independent. $\Psi_{x}(k)$ is doubly periodic with respect to $k$:
$\Psi_{x}(k+2w_{i})=\Psi_{x}(k),\quad i=1,2.$ (2.8)
@@ 2.2 Connection with the Marchenko equation
In the following we consider the case where the moduli $g_{2}$ and $g_{3}$ are real and the discriminant $\Delta$ is positive. In this case, according to Proposition A.2, the half-periods $w_{1}$ and $w_{2}$ are real and purely imaginary, respectively. We also suppose $\mathrm{Im}w_{2}>0$ without loss of generality.
Note that $\sigma(z)$ takes zero when $z=2mw_{1}+2nw_{2}$, $m,n\in\mathbb{Z}$, and even when $z\in\mathbb{R}$ there are periodic zeros when $z=2mw_{1}$ for $m\in\mathbb{Z}$, which will bring singularities to elliptic solitons. In order to get real-valued non-singular solutions, we shift the zeros from the real axis to the complex plane by performing a transformation $x\mapsto x+w_{2}$. Thus we introduce
$\widetilde{\Psi}_{x}(k)$ $\doteq\frac{\sigma(x+k+w_{2})}{\sigma(x+w_{2})\sigma(k)}e^{-\zeta(k)x-\zeta(w_{2})k},$ (2.25a)
$\widetilde{\Psi}_{x}(k,l)$ $\doteq\frac{\sigma(x+k+l+w_{2})}{\sigma(x+w_{2})\sigma(k+l)}e^{-(\zeta(k)+\zeta(l))x-\zeta(w_{2})(k+l)}.$ (2.25b)
@@
$v(x;a,b)=\frac{\sigma(x+w_{2}+a+b)}{\sigma(x+w_{2})\sigma(a+b)}e^{-\zeta(w_{2})(a+b)-\frac{(a+b)\zeta(w_{1})}{w_{1}}x},$ (2.26b)
$p(a)=a\zeta(w_{1})-w_{1}\zeta(a).$ (2.26c)
@@
$\widetilde{\Psi}_{x}(k+2w_{1},l)=-\widetilde{\Psi}_{x}(k,l),$ (2.27a)
$\widetilde{\Psi}_{x}(k+2w_{2},l)=\widetilde{\Psi}_{x}(k,l).$ (2.27b)
@@
$p(k+2w_{1})=p(k),\quad p(k+2w_{2})=p(k)+\pi\mathrm{i},\quad p(-k)=-p(k).$ (2.29)
When $k\in(0,w_{2}]$ and $(w_{1},w_{3}]$, $p(k)$ is purely imaginary; when $k\in(w_{2},w_{3})$, $p(k)$ is complex, and the imaginary part is always $\frac{\pi}{2}$; when $k\in(0,w_{1}]$, $p(k)$ is real, it varies monotonically from $\infty$- ∞ at the origin to 00 at $w_{1}$.
@@
(1). The proof of (2.27a) employs formula $w_{2}\zeta(w_{1})-w_{1}\zeta(w_{2})=\frac{\pi\mathrm{i}}{2}$ and the quasi-periodicity properties (2.7). The proof of (2.27b) is straight forward.
@@
$u(x,y,t)=-2\wp(x+w_{2})-2\partial_{x}\iint_{D}\psi(x,y,t;l)\gamma_{l^{\prime}}(y,t)\widetilde{\Psi}_{x}(l^{\prime})\mathrm{d}\lambda(l,l^{\prime}).$ (2.31)
@@
In what follows, we restrict ourselves to the case where $x,g_{2},g_{3}\in\mathbb{R}$, $w_{1}>0$, $\mathrm{Im}w_{2}>0$ and $\Delta>0$. Particularly, we take $k,l^{\prime}\in(0,w_{1})$. It then follows from item (2) and (3) of Proposition 2.2 that $v(x;k,l^{\prime})$ is bounded on the real axis, and $p(k)+p(l^{\prime})<0$, and consequently, $\widetilde{\Psi}_{x}(k,l^{\prime})$ exponentially decreases to zero as $x\to+\infty$. Thus, from (2.34) we have
@@ 2.3 Elliptic soliton solutions
where $B_{i,j}\in\mathbb{C}$, $k_{i}$ and $k^{\prime}_{j}$ belong to the fundamental periodic parallelogram formed by $2w_{1}$ and $2w_{2}$, and we assume $w_{1}>0,\mathrm{Im}w_{2}>0$. Then the integral equation (2.9) reduces to a linear algebraic system
@@
$\tau_{N}=\sum_{J\subset S}\left(\prod_{i\in J}c_{i}\right)\!\!\left(\mathop{\rm{\prod}}_{i<j\in J}A_{i,j}\right)\!\frac{\sigma(x+w_{2}+\sum_{i\in J}(k_{i}+k^{\prime}_{i}))}{\sigma(x+w_{2})\prod_{i\in J}\sigma(k_{i}+k^{\prime}_{i})}\mathrm{exp}\left(\sum_{i\in J}\theta_{[e]}(x,y,t;k_{i},k^{\prime}_{i})\right),$ (2.52)
@@
$c_{i}=\widetilde{c}_{i}\mathrm{\exp}\left(-\zeta(w_{2})(k_{i}+k^{\prime}_{i})\right),~{}~{}~{}\widetilde{c}_{i}\in\mathbb{R}.$ (2.53)
@@
$\tau_{1}=1+c_{1}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+w_{2})}{\sigma(x+w_{2})\sigma(k_{1}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})},$ (2.54b)
@@
$\tau_{2}=$ $1+c_{1}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+w_{2})}{\sigma(x+w_{2})\sigma(k_{1}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})}+c_{2}\frac{\sigma(x+k_{2}+k^{\prime}_{2}+w_{2})}{\sigma(x+w_{2})\sigma(k_{2}+k^{\prime}_{2})}e^{\theta_{[e]}(x,y,t;k_{2},k^{\prime}_{2})}$
$+c_{1}c_{2}A_{1,2}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+k_{2}+k^{\prime}_{2}+w_{2})}{\sigma(x+w_{2})\sigma(k_{1}+k^{\prime}_{1})\sigma(k_{2}+k^{\prime}_{2})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})+\theta_{[e]}(x,y,t;k_{2},k^{\prime}_{2})}.$ (2.55b)
@@
[…]
Then $u(x,y,t)$ is a real function when $x,y,t,k_{j},l_{j},\widetilde{c}_{j}\in\mathbb{R}$. Dynamics of these solutions are depicted in Fig.1, which describe line solitons on a periodic background that is due to $-2\wp(x+w_{2})$.
@@
$\tau_{2}=1+c_{1}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+w_{2})}{\sigma(x+w_{2})\sigma(k_{1}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})}+c_{2}\frac{\sigma(x+k_{2}+k^{\prime}_{1}+w_{2})}{\sigma(x+w_{2})\sigma(k_{2}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{2},k^{\prime}_{1})}.$ (2.57)

### v3

@@ 1 Introduction
A quite recent progress of the DL approach is the establishment of its elliptic scheme involving infinite matrix structure and ‘discrete’ Lamé-type PWFs [44]. The Lamé function is a solution of the Lamé equation (known as the Schrödinger equation with an elliptic potential) $y^{\prime\prime}+(A+B\wp(x))y=0$ when taking $A=-\wp(k)$ and B=2B=-2. The solution reads $y=\frac{\sigma(x+k)}{\sigma(x)\sigma(k)}e^{-\zeta(k)x}$, where $\sigma,\zeta$ (and $\wp$) are the Weierstrass functions. For the KdV equation, there are exact solutions involving the Lamé function, which emerged in [30] and [66] 1970s. In 2010, Nijhoff and Atkinson [38] employed a discrete analogue of the Lamé function and constructed solutions for the Adler-Bobenko-Suris [3] (except Q4 equation) lattice equations. The obtained solutions are presented in terms of the elliptic Cauchy matrix and called ‘elliptic solitons’ to distinguish from the algebro-geometric (finite-gap) solutions (cf.[10]). Later, their treatment was extended to the lattice KP equations [69]. The elliptic DL scheme established in [44] is based on the infinite matrix structures. The concept of ‘elliptic $N$th root of unity’ was introduced to define discrete dispersion relations as well as make reductions from the lattice KP system to the lower dimensional equations, namely, the discrete KdV and Boussinesq equations. As for the elliptic solitons for the continuous equations, a recent progress is due to [32], which established a framework of bilinear approach that involves the Lamé function, including formulae of $\tau$ functions, vertex operators and bilinear identities for the KP hierarchy and their reductions. Periodic degenerations (i.e. trigonometric/hyperbolic and rational forms) of the elliptic matter were also discussed in [32]. Later, the study of [32] was extended to discrete case in [33]. More recently, the consideration of such type of elliptic soliton solutions from Sato’s approach was developed by Kakei [27], where the Darboux transformation based on pseudo-differential operators and the Lamé function was established getting elliptic $N$-soliton solutions for the whole KP hierarchy and the KdV hierarchy. The elliptic solitons are also related to a recent work by Nakayashiki [36].
@@
In this paper, we will extend the Fokas-Ablowitz DL scheme to the elliptic soliton case222 In the paper by elliptic soliton we mean the solutions involving the Lamé function instead of the usual exponential functions for the ordinary solitons, which follows the naming in [38]. Strictly speaking, “elliptic solitons” should be elliptic in the spatial variable $x$ (cf.[28]). We will provide a brief review on such “elliptic solitons” in Appendix C and formulate a mechanism to obtain non-singular real-valued elliptic soliton solutions. First, for the KP equation, we will present an integral equation involving elliptic PWFs and Lamé function, using which we can define a wave function and solution for the KP equation. This DL scheme will be confirmed by checking the wave function and solution satisfy the KP Lax pair. Real-valuedness of the Weierstrass functions will be discussed (see Appendix A). In light of this, we are able to construct a Marchenko equation for elliptic soliton solutions of the KP equation. By taking special measure in the integral equation, we may obtain explicit expressions for the elliptic solitons (as well as $\tau$ functions) of the KP equation. Some solutions will be illustrated. In principle, employing the notion of elliptic $N$th root of unity introduced in [44], we can reduce the elliptic DL scheme, Marchenko equation and solutions of the KP equation to those of the KdV equation and the Boussinesq equation. However, in practice, there are many details to be fixed, which will be elaborated in the paper.
@@
We organize the paper as follows. In section 2, we set up the elliptic DL scheme for the KP equation, construct its Marchenko equation and derive formulae of the elliptic $N$-soliton solution and $\tau$-function. Then, in section 3, we elaborate reduction details so that the elliptic DL schemes, Marchenko equations and elliptic solitons can be constructed for the KdV equation and Boussinesq equation. Conclusions are given in section 4. There are three Appendix sections. The first one is devoted to the real-valuedness of the Weierstrass functions. The second one consists of the analysis about elliptic cube roots. The third one includes a brief review on “elliptic solitons”.
@@ 2.1 Elliptic DL scheme
and the KP equation arises from the compatibility [P,M]=PMMP=0[P,M]=PM-MP=0. When $u(x,y,t)=u_{0}(x)$, the above Lax pair admits a special solution
@@
$\wp(z+2mw+2nw^{\prime})=\wp(z),~~m,n\in\mathbb{Z}$
where $w$ and $w^{\prime}$ are the fundamental half-periods, $\zeta(z)$ and $\sigma(z)$ are odd and quasi-periodic, in the sense that
$\zeta(z+2mw+2nw^{\prime})=\zeta(z)+2m\zeta(w)+2n\zeta(w^{\prime}),$ (2.7a)
$\sigma(z+2mw+2nw^{\prime})=(-1)^{m+n+mn}\sigma(z)e^{(z+mw+nw^{\prime})(2m\zeta(w)+2n\zeta(w^{\prime}))}.$ (2.7b)
When $k\neq w,w^{\prime}$, $\Psi_{x}(k)$ and $\Psi_{x}(-k)$ are linearly independent. $\Psi_{x}(k)$ is doubly periodic with respect to $k$:
$\Psi_{x}(k+2mw+2nw^{\prime})=\Psi_{x}(k).$ (2.8)
@@ 2.2 Connection with the Marchenko equation
In the following we consider the case where the moduli $g_{2}$ and $g_{3}$ are real and the discriminant $\Delta$ is positive. In this case, according to Proposition A.2, the half-periods $w$ and $w^{\prime}$ are real and purely imaginary, respectively. We also suppose $\mathrm{Im}w^{\prime}>0$ without loss of generality.
Note that $\sigma(z)$ takes zero when $z=2mw+2nw^{\prime}$, $m,n\in\mathbb{Z}$, and even when $z\in\mathbb{R}$ there are periodic zeros when $z=2mw$ for $m\in\mathbb{Z}$, which will bring singularities to elliptic solitons. In order to get real-valued non-singular solutions, we shift the zeros from the real axis to the complex plane by performing a transformation $x\mapsto x+w^{\prime}$. Thus we introduce
$\widetilde{\Psi}_{x}(k)$ $\doteq\frac{\sigma(x+k+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k)}e^{-\zeta(k)x-\zeta(w^{\prime})k},$ (2.25a)
$\widetilde{\Psi}_{x}(k,l)$ $\doteq\frac{\sigma(x+k+l+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k+l)}e^{-(\zeta(k)+\zeta(l))x-\zeta(w^{\prime})(k+l)}.$ (2.25b)
@@
$v(x;a,b)=\frac{\sigma(x+w^{\prime}+a+b)}{\sigma(x+w^{\prime})\sigma(a+b)}e^{-\zeta(w^{\prime})(a+b)-\frac{(a+b)\zeta(w)}{w}x},$ (2.26b)
$p(a)=a\zeta(w)-w\zeta(a).$ (2.26c)
@@
$\widetilde{\Psi}_{x}(k+2w,l)=-\widetilde{\Psi}_{x}(k,l),$ (2.27a)
$\widetilde{\Psi}_{x}(k+2w^{\prime},l)=\widetilde{\Psi}_{x}(k,l).$ (2.27b)
@@
$p(k+2w)=p(k),\quad p(k+2w^{\prime})=p(k)+\pi\mathrm{i},\quad p(-k)=-p(k).$ (2.29)
When $k\in(0,w^{\prime}]$ and $(w,w+w^{\prime}]$, $p(k)$ is purely imaginary; when $k\in(w^{\prime},w+w^{\prime})$, $p(k)$ is complex, and the imaginary part is always $\frac{\pi}{2}$; when $k\in(0,w]$, $p(k)$ is real, it varies monotonically from -$\infty$ at the origin to 0 at $w$.
@@
(1). The proof of (2.27a) employs formula $w^{\prime}\zeta(w)-w\zeta(w^{\prime})=\frac{\pi\mathrm{i}}{2}$ and the quasi-periodicity properties (2.7). The proof of (2.27b) is straight forward.
@@
$u(x,y,t)=-2\wp(x+w^{\prime})-2\partial_{x}\iint_{D}\psi(x,y,t;l)\gamma_{l^{\prime}}(y,t)\widetilde{\Psi}_{x}(l^{\prime})\mathrm{d}\lambda(l,l^{\prime}).$ (2.31)
@@
In what follows, we restrict ourselves to the case where $x,g_{2},g_{3}\in\mathbb{R}$, $w>0$, $\mathrm{Im}w^{\prime}>0$ and $\Delta>0$. Particularly, we take $k,l^{\prime}\in(0,w)$. It then follows from item (2) and (3) of Proposition 2.2 that $v(x;k,l^{\prime})$ is bounded on the real axis, and $p(k)+p(l^{\prime})<0$, and consequently, $\widetilde{\Psi}_{x}(k,l^{\prime})$ exponentially decreases to zero as $x\to+\infty$. Thus, from (2.34) we have
@@ 2.3 Elliptic soliton solutions
where $B_{i,j}\in\mathbb{C}$, $k_{i}$ and $k^{\prime}_{j}$ belong to the fundamental periodic parallelogram formed by $2w$ and $2w^{\prime}$, and we assume $w>0,\mathrm{Im}w^{\prime}>0$. Then the integral equation (2.9) reduces to a linear algebraic system
@@
$\tau_{N}=\sum_{J\subset S}\left(\prod_{i\in J}c_{i}\right)\!\!\left(\mathop{\rm{\prod}}_{i<j\in J}A_{i,j}\right)\!\frac{\sigma(x+w^{\prime}+\sum_{i\in J}(k_{i}+k^{\prime}_{i}))}{\sigma(x+w^{\prime})\prod_{i\in J}\sigma(k_{i}+k^{\prime}_{i})}\mathrm{exp}\left(\sum_{i\in J}\theta_{[e]}(x,y,t;k_{i},k^{\prime}_{i})\right),$ (2.52)
@@
$c_{i}=\widetilde{c}_{i}\mathrm{\exp}\left(-\zeta(w^{\prime})(k_{i}+k^{\prime}_{i})\right),~~~\widetilde{c}_{i}\in\mathbb{R}.$ (2.53)
@@
$\tau_{1}=1+c_{1}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k_{1}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})},$ (2.54b)
@@
$\tau_{2}=$ $1+c_{1}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k_{1}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})}+c_{2}\frac{\sigma(x+k_{2}+k^{\prime}_{2}+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k_{2}+k^{\prime}_{2})}e^{\theta_{[e]}(x,y,t;k_{2},k^{\prime}_{2})}$
$+c_{1}c_{2}A_{1,2}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+k_{2}+k^{\prime}_{2}+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k_{1}+k^{\prime}_{1})\sigma(k_{2}+k^{\prime}_{2})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})+\theta_{[e]}(x,y,t;k_{2},k^{\prime}_{2})}.$ (2.55b)
@@
$w=1.22569...,\qquad w^{\prime}=\mathrm{i}\,1.49673.$... (2.56)
Then $u(x,y,t)$ is a real function when $x,y,t,k_{j},l_{j},\widetilde{c}_{j}\in\mathbb{R}$. Dynamics of these solutions are depicted in Fig.1, which describe line solitons on a periodic background that is due to $2\wp(x+w^{\prime})$.
@@
$\tau_{2}=1+c_{1}\frac{\sigma(x+k_{1}+k^{\prime}_{1}+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k_{1}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{1},k^{\prime}_{1})}+c_{2}\frac{\sigma(x+k_{2}+k^{\prime}_{1}+w^{\prime})}{\sigma(x+w^{\prime})\sigma(k_{2}+k^{\prime}_{1})}e^{\theta_{[e]}(x,y,t;k_{2},k^{\prime}_{1})}.$ (2.57)